Session: May 1998 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Westerbork (Tony Foley at Wed Jun 10 11:58:05 METDST 1998) Success
No known problems

Torun (K. Borkowski at Wed Jun 10 12:01:03 METDST 1998) Success
No known problems.

Noto (C.Trigilio at Wed Jun 10 12:19:12 METDST 1998) Success

Jodrell (Alastair Gunn at Wed Jun 10 12:28:16 METDST 1998) Success
No known problems.

Medicina (T.Venturi at Wed Jun 10 12:48:09 METDST 1998) Minor failures
Problems with the capstan. Tape flying over capstan. One scan lost (from 9.13 to 9.38 ut) and others may have been affected.

Calibration (Jean\-Francois Desmurs at Thu Jun 11 12:43:03 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the mai98 directory under the name v090a2.antab. Read the comments in the file to get further information.

Onsala (A. Polatidis at Thu Jun 11 12:55:56 METDST 1998) Failure
The experiment was not observed due to high winds and a problem with the declination brake of the 25m antenna. The brake was fixed on Thursday afternoon.

Effelsberg (dgraham at Fri Jun 12 15:07:03 METDST 1998) Minor failures
Antenna drive problems around 2130,0030,0330, each time about 30 minutes lost.